OpenStudy (anonymous):

How many electrons should be removed from each of the two similar spheres each of 10g so that electrostatic repulsion be balanced by gravitational force.

OpenStudy (anonymous):

I think that you mean removing some electrons and adding the to the other ball! It's obvious that the forces should be equal,(m1=m2=m, q1=q2=q) F1=F2=>Gm1*m2/r^2 = kq1*q2/r^2 => q^2=G*m^2/k G=6.67300 × 10-11 k=9.109*10^9 m=.01kg =>q=2.723*10^-12 q=ne=> n = 1.7018*\[10^{7}\]